OpenStudy (anonymous):

5. Find the solution of y = –4x + 9 for x = 5. A) (5, 1) B) (5, –11) C) (5, –56) D) (29, 5)

OpenStudy (anonymous):

x is 5 y is -11

OpenStudy (wolfboy):

Yes because

OpenStudy (anonymous):

plug x into the eqn fr y

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = -4(5) + 9}\) This is it :D

OpenStudy (diyadiya):

Plug in the value of x y= -4(5) + 9 and find y and your answer is in the form (x,y)

OpenStudy (anonymous):

so...b?

OpenStudy (wolfboy):

5*-4=-20 -20+p=-11

Parth (parthkohli):

(x,y) ---> x is already defined as 5.

OpenStudy (wolfboy):

*9 not p

OpenStudy (wolfboy):

so its x=5y=-11

OpenStudy (wolfboy):

So its B

OpenStudy (anonymous):

thx for your help :)

OpenStudy (wolfboy):

np thats what we are here for

OpenStudy (anonymous):

also...wolfy nice pic LOL

OpenStudy (wolfboy):

thx

Parth (parthkohli):

Giving the answer....not always the best idea ever

OpenStudy (wolfboy):

I showed her every step

OpenStudy (anonymous):

srry :(

OpenStudy (wolfboy):

I did not immediately give the answer

OpenStudy (anonymous):

although it should be fine if you start with the explanation then give the answer

OpenStudy (wolfboy):

Same difference

OpenStudy (anonymous):

LOL true

Parth (parthkohli):

Not always...read the code of conduct

Parth (parthkohli):

nvm

OpenStudy (anonymous):

im leavign dont wanna b in a fight

OpenStudy (wolfboy):

I was gonna do that but the other person blurted out the answer

OpenStudy (anonymous):

LOL

razor99 (razor99):

good idea

OpenStudy (wolfboy):

So i showed her how that is all

razor99 (razor99):

@TaylorBaugher2